# Secret Breach Detection in Source Code with Large Language Models

### v1

@@ V Threats to Validity
Conclusion validity reflects the accuracy of the interpretations made from our experimental results. To ensure robust conclusions, we used standard evaluation metrics—precision, recall, and F1-score—along with confusion matrix analysis. Still, model performance can vary across deployment scenarios, especially in noisy or highly imbalanced environments. We attempted to reduce individual evaluation bias by relying on automated classification and consistent test datasets. However, model-specific peculiarities or overfitting to the fine-tuning dataset could skew the results. Future replication studies can help verify our findings across alternate settings.
VI Conclusion
@@
Looking forward, our approach can be extended by integrating it into CI/CD pipelines and IDEs for real-time secret leak prevention. Future work may also explore ensemble methods or uncertainty-based rejection mechanisms to further minimize or eliminate false negatives. Additionally, expanding detection to other artifacts such as issue reports and improving generalization through domain adaptation for proprietary and multilingual codebases remains an important direction. Furthermore, the development of more effective and adaptive prompting techniques for security tasks may further boost the performance of pre-trained models.

### v2

@@ VI Threats to Validity
Conclusion validity reflects the accuracy of the interpretations made from our experimental results. To ensure robust conclusions, we used standard evaluation metrics—precision, recall, F1 and F2-score along with confusion matrix analysis. Still, model performance can vary across deployment scenarios, especially in noisy or highly imbalanced environments. We attempted to reduce individual evaluation bias by relying on automated classification and consistent test datasets. However, model-specific peculiarities or overfitting to the fine-tuning dataset could skew the results. Future replication studies can help verify our findings across alternate settings.
VII Conclusion
@@
Looking ahead, we plan to integrate our approach into CI/CD pipelines and IDEs for real-time secret detection. Future work includes exploring ensemble models, uncertainty-based rejection to reduce false negatives, and extending detection to other artifacts. We also aim to enhance generalization via domain adaptation and improve adaptive prompting for security-focused tasks.